OpenStudy (anonymous):

There are two ponds of the same size. The first one has 1 water lily and the second has 2 lilies. Every 5 seconds, each lily in both ponds doubles itself. The first pond takes 5 minutes to fill. How much time will the second pond need?

OpenStudy (anonymous):

Answer is: 4 minutes 55 seconds, but please explain the algebra.

OpenStudy (anonymous):

okayy.. we know each 5 sec lily doubles.. 5 min 300 secs in this 300 secs the lily doubles itself 60 times.. the starting term is 1 so the seq goes on.. now what is this seq.. its just 2^n. 2^0=1 2^1=2 2^2=4 so on.. now 2^ 59=5.7 * 10^17 would be the amount of lilies required to fill the pond assuming both ponds to be same in size.. when the number of lilies become 2^59 in pond 2.. its full pond two starts with 2 lilies ie 2^1.. this means that its 1 term ahead of pond 1.. thus pond 2 blossoms 5 secs before pond 1 :)